# Minimalistic and Scalable Quantum Reservoir Computing Enhanced with Feedback

### v1

@@ II.2 Mackey-Glass Task
For CRC based on classical echo state networks (ESN), performance enhancement resulting from the increased number of neurons is plotted by the red solid lines in Figs. 3(a)(b), with details discussed in the "Supplemental Information". The advantage of QRC over CRC lies in the exponentially increased number of quantum basis states that span a large QRC Hilbert space, which underlies a relatively small number of measured neurons.
@@
The effect of solely increasing the dimension of the QRC Hilbert space is demonstrated in Fig. 3(c), where the number of measured readouts is fixed at $4$ (including $2$ from the cavity and $2$ from an atom), while additional unmeasured atoms are added for the only purpose of increasing the number of basis states. Performance improvements are observed in both the no-feedback and 2-feedback cases, indicating the inherent connections and collaborations between the measured atom and the added atoms. The effect tends to saturate in the 4-feedback case as the use of all the $4$ available feedback channels tends to dominate the performance. The similar effect has been observed for Ising models of QRC Fujii2017 ; chen2019learning . A reason for this effect could be that, for the particular tasks considered, more complex fading memory maps generated by the QRC as the Hilbert space is increased are able to better capture features in the task to be learned. However, the improvement is expected to eventually plateau for a high enough dimension of the Hilbert space. In Fig. 3(c), we also compare the results with the classical ESN. One of the most striking features of QRC over CRC is that increasing the number of unmeasured neurons improves the performance in QRC whereas the CRC does not improve the performance even if we increase the number of unmeasured neurons. We discuss this further in the "Supplemental Information".
@@
Figure 4: Performance testing for the Mackey-Glass task using one-atom QRC with various $Delay$ and decay rates $\kappa$. (a) NRMSE as a function of $Delay$, with no feedback or 4 feedbacks, for $\kappa=10$. (b)(c) The actual (red) and target (blue) outputs for $Delay=200$ and $Delay=2$, corresponding to the points marked by letters "b" and "c" in panel (a), respectively. (d) NRMSE as a function of $\kappa$, with no feedback or 4 feedbacks, for $Delay=20$. (e)(f) The actual (red) and target (blue) outputs for $\kappa=10^{5}$ and $\kappa=10$, corresponding to the points marked by letters "e" and "f" in panel (d), respectively. Parameters: $\omega_{1}=20$, $g_{1}=30$, $\omega_{c}=40$, $\epsilon=20$.
The impacts of $Delay$ and the decay rate $\kappa$ are shown in Fig. 4. In the Mackey-Glass task, a longer $Delay$ necessitates a stronger memory for the reservoir to "remember" more past information from the input in order to forecast the future output, making the task increasingly challenging. The decay rate $\kappa$ characterizes the loss of quantum memory caused by quantum dissipation, as define in the stochastic master equation introduced in the "Methods" section. A larger $\kappa$ tends to cause the reservoir to "forget" input information more quickly due to photon leakage and atomic spontaneous emission. However, a larger $\kappa$ is beneficial for stronger measurements with reduced uncertainties in readouts. Therefore, selecting a balanced $\kappa$ could be crucial in practice.
@@ II.3 Sine-square Waveform Classification Task
The objective of the classification task is to determine whether each input data point belongs to a sine or a square waveform. The time-dependent input, $f_{k}$, comprises $110$ randomly generated sine and square waveforms. Among these, $10$ waveforms are allocated for memory fading, $50$ for training, and $50$ for testing. Each waveform is discretized into $N_{ss}$ points, leading to a time step of $\Delta t=2\pi/\left(N_{ss}\omega_{ss}\right)$, where $\omega_{ss}$ denotes the oscillation frequency of the input. Figure 5(a) depicts these discretized time points (red dots) across the first $10$ waveforms during the testing phase. The corresponding input modified by feedback is plotted in Fig. 5(b). The target output, $\bar{y}_{k}$, aimed at classifying the input signal, is set to 00 if the input point belongs to a square waveform, and $1$ if it belongs to a sine waveform.
@@
The sine-square waveform classification is a nonlinearity task that requires the reservoir to process a linearly inseparable input dataset containing abrupt changes. In a linear, closed quantum system, the smoothly evolving expectation values in readouts, $\left\langle c\right\rangle\propto\exp\left(-i\omega_{c}t\right)$ and $\left\langle\sigma_{i}\right\rangle\propto\exp\left(-i\omega_{i}t\right)$, are unable to capture these high-frequency, abrupt shifts. The introduction of nonlinearity, originating from the coherent laser driving and decay in the open quantum system, enhances the reservoir’s capability to promptly respond to the abrupt changes in the input. This is evidenced by the distinct readout measurements in Fig. 5(c)-(f), which correspond one-on-one to the input waveforms in Fig. 5(a).
Figure 6: Performance testing for the sine-square waveform classification task with various reservoir scales. (a)(b) Tested NRMSE plotted against the number of atoms (or neurons), using all available readout channels, 00, $2$, or $4$ feedback channels, and regular linear regression (solid lines) or polynomial regression (dashed lines), with a comparison to CRC employing classical ESN (red solid line), where the parameters $\omega_{i}$ and $g_{i}$ align with those in Fig. 3(a)(b). (c)(d) Actual (red) and target (blue) outputs from QRC utilizing one atom with no feedback with linear regression, and three atoms with $4$ feedbacks with polynomial regression, corresponding to the points marked by "c" and "d" in panel (b), respectively. Parameters: $\omega_{c}=40$, $\kappa=10$, $\epsilon=20$, $\omega_{ss}=10$, and $N_{ss}=16$.
@@
The performance enhancements resulted from the increased number of atoms, feedback mechanism, and polynomial regression, along with the comparison between QRC and CRC, are shown in Figs. 6(a)(b). It is observed that the performance associated with polynomial regression, represented by the dashed lines, tends to saturate at $4$ atoms. This performance saturation suggests that the maximum performance achievable with the current sample size, associated with $N_{ss}=16$, has been attained. The enhanced performance is illustrated in Fig. 6(d), where discrepancies between the actual and target outputs primarily occur at the first point in each waveform. These discrepancies characterize the short-time responses of the quantum reservoir to the abrupt shifts in the input waveform. The dependence of QRC performance on various samples sizes $N_{ss}$ and the influence of abrupt shifts in the input are detailed in the "Supplemental Information".
@@ III Discussion
We have introduced a paradigm for minimalistic quantum reservoirs that enhances quantum memory and nonlinear data processing capabilities. This approach incorporates a feedback mechanism that modifies the inputs based on observable readouts. Specifically, the inputs are encoded into the intensity of the cavity-driving laser, enabling a practical implementation of feedback by adjusting the laser’s intensity. To further augment the system, we employ polynomial regression to append quadratic combinations of observable expectations to the readouts, introducing an additional layer of complexity and capability.
@@
These novel components significantly enhance the performance of QRC for small reservoirs containing just one to five atoms. To obtain the readouts in practice, we propose the use of continuous quantum measurements via homodyne detection of cavity quadratures and atomic spins. This approach allows for the simultaneous measurement of non-commuting observables Wei2008 ; Ruskov2010 ; Hacohen2016 ; Ochoa2018 , avoiding the need for tomography and making it more practical than the measurement of probability distributions in quantum basis states suggested in prior works Fujii2017 ; Ghosh2019 ; Dudas2023 ; Angelatos2021 .
@@
Our proposed quantum optical reservoir offers convenient scalability compared to reservoirs built upon quantum networks like the Ising Fujii2017 and Fermi-Hubbard Ghosh2019 models. This advantage arises from the practical coupling between atoms and a single-mode cavity field in our system. Specifically, to scale up the reservoir, we simply need to add more atoms into the cavity. These atoms naturally couple with the cavity field, which in turn automatically couples them with the rest of the atoms in the reservoir.
@@
Moreover, the number of quantum basis states in our reservoir scales exponentially with the number of atoms, following $2^{N_{\text{atom}}}$. This leads to faster growth compared to reservoirs not based on quantum networks, such as a single Kerr nonlinear oscillator Govia2021 or two coupled linear oscillators Dudas2023 . This exponential scaling has been crucial in demonstrating the advantages of QRC over CRC, as shown in Figs 3(a)(b) and 6(a)(b).
@@ IV.1 Continuous quantum measurement for observable readouts
Continuous quantum measurements are simulated for both the cavity field and individual atoms. For the cavity field, homodyne detection of two orthogonal quadratures involves splitting the system’s output beam into two using a beam-splitter, followed by homodyning each beam with the same local oscillator, differing by a phase shift of $\pi/2$ Wiseman2010 . Similarly, for atoms, homodyne detection is performed for spontaneous emissions Wiseman2001 . These measurements can be conducted concurrently Wei2008 ; Ruskov2010 ; Hacohen2016 ; Ochoa2018 . Consequently, the continuous measurement process is described by the stochastic master equation VPB79 ; VPB88 ; VPB91a ; HC93 ; Wiseman2010 ; BvHJ07 ; Nurdin14
@@
[…]
@@
[…]
@@
for any stochastic collapse operator $a$. The continuous quantum measurements of the observables $Q$, $P$, $\sigma_{x,i}$, and $\sigma_{y,i}$ are associated with the stochastic collapse operators $\sqrt{\kappa_{c}}c$, $i\sqrt{\kappa_{c}}c$, $\sqrt{\kappa_{i}}\sigma_{i}$, and $i\sqrt{\kappa_{i}}\sigma_{i}$, respectively Wiseman2010 ; BvHJ07 ; Nurdin14 . The randomness of the measurement records is taken into account by the Wiener increments, $dW_{Q\left(P\right)}$ and $dW_{x\left(y\right),i}$, each of which selects a random number from a Gaussian probability distribution with a width of $dt$. The efficiencies of various detection channels are incorporated into the Wiener increments. Each measurement detects the continuous currents in cavity and atom channels with noises, with the measurement records given by $\left\langle Q\left(P\right)\right\rangle_{J}+dW_{Q\left(P\right)}/dt$ and $\left\langle\sigma_{x\left(y\right),i}\right\rangle_{J}+dW_{x\left(y\right),i}/dt$, where the expectation values are computed using $\rho_{J}$ from Eq. (10).
@@
[…]
@@
accompanied by the averaged measurement records $\left\langle Q\left(P\right)\right\rangle$ and $\left\langle\sigma_{x\left(y\right),i}\right\rangle$, where the expectation values are computed using $\rho$ from Eq. (13). To maintain a constant total decay rate $\kappa$ as $N_{atom}$ increases, the decay rate of the cavity or each atom is assumed to be $\kappa_{c\left(i\right)}=\kappa/\left(2N_{atom}+2\right)$. A higher $\kappa$ reduces the uncertainty in the readout measurement, commonly referred to a strong measurement, while a lower $\kappa$ corresponds to a weak measurement Fuchs2001 . The influence of the magnitude of $\kappa$ on QRC performance is demonstrated in Fig. 4.
To investigate the dynamics of QRC under the feedback scheme from the time $t_{k}$ to $t_{k+1}$, the quantum reservoir is initialized with the modified input $\widetilde{f}_{k}$ and $\rho\left(t_{k}\right)$ at the time $t_{k}$. Here, $\widetilde{f}_{k}$ is determined by the original input $f_{k}$ and the expectation values calculated using $\rho\left(t_{k}\right)$. The evolution from $t_{k}$ to $t_{k+1}$ simulated by the master equation gives rise to $\rho\left(t_{k+1}\right)$, which are then utilized to compute $\widetilde{f}_{k+1}$ and initialize the next feedback cycle. From the time $t_{k}$ to $t_{k+1}$, the "mesolve" and "smesolve" methods in the QuTiP library are employed to solve the stochastic master equation (10) and the deterministic master equation (13), respectively. In our setup, the two master equations lead to the roughly same readouts $x_{nk}$ when the number of averaged stochastic measurement trajectories in the "smesolve" method reaches $5000$.

### v2

@@ II.2 Mackey-Glass Task
For CRC based on classical echo state networks (ESN), performance enhancement resulting from the increased number of neurons is plotted by the red solid lines in Figs. 3(a)(b), with details discussed in the “Supplementary Information”. The primary advantage of QRC over CRC lies in the entanglement of atomic states through coupling with a shared cavity field. The exponentially increased number of atomic basis states achieves a high-dimensional Hilbert space for QRC, despite only a relatively small number of observable readouts are being measured.
@@
The effect of solely increasing the dimension of the QRC Hilbert space is demonstrated in Fig. 3(c), where the number of measured readouts is fixed at $4$ (including $2$ from the cavity and $2$ from an atom), while additional unmeasured atoms are added for the only purpose of increasing the number of basis states. Performance improvements are observed in both the no-feedback and 2-feedback cases, indicating the inherent connections and collaborations between the measured atom and the added atoms. The effect tends to saturate in the 4-feedback case as the use of all the $4$ available feedback channels tends to dominate the performance. The similar effect has been observed for Ising models of QRC [Fujii2017, chen2019learning]. A reason for this effect could be that, for the particular tasks considered, more complex fading memory maps generated by the QRC as the Hilbert space is increased are able to better capture features in the task to be learned. However, the improvement is expected to eventually plateau for a high enough dimension of the Hilbert space. In Fig. 3(c), we also compare the results with the classical ESN. One of the most striking features of QRC over CRC is that increasing the number of unmeasured neurons improves the performance in QRC (black and blue solid lines in Fig. 3(c)), whereas the CRC does not improve the performance even if we increase the number of unmeasured neurons (flat red line in Fig. 3(c)). In the “Supplementary Information”, the performance of CRC is evaluated by fixing the measured neurons while increasing the number of unmeasured neurons, and no performance improvement is observed for CRC.
@@
The feedback parameters $V_{n_{f}}$ are optimized using three global optimization methods: differential evolution, brute force, and brute force combined with the Nelder-Mead algorithm. The corresponding test results are shown in Fig. 3(h), with further details provided in the “Methods” section. Our results suggest that the feedback mechanism yields a noticeable improvement in QRC performance, even when only a small number of $V_{n_{f}}$ values are explored within a constrained parameter space and only a single feedback channel is activated. Thus, rigorous global optimization of $V_{n_{f}}$ is not necessary to achieve substantial performance enhancement. For example, in the brute force method, the $V_{n_{f}}$ space is discretized with a step size of $0.5$ and the search boundaries are limited to $V_{n_{f}}\in\left[-3,3\right]$. Referring to these settings and the 1-feedback case shown in Fig. 3(h), the NRMSE improves from approximately $0.11$ to $0.09$ by evaluating only $13$ different values of $V_{n_{f}}$. The NRMSE profile as a function of a single feedback parameter $V_{m}$–with all other feedback channels deactivated (i.e., $V_{n_{f}\neq m}=0$)–is presented in Fig. S1(a-d) of the “Supplementary Information.” These plots demonstrate performance enhancement resulted from various options for $V_{m}$. Additionally, as shown in Fig. S1, high-frequency oscillations in the readouts, related to a limit-cycle-like behavior, are observed in certain parameter regions.
Figure 4: Performance testing for the Mackey-Glass task using one-atom QRC with various DelayDelay and decay rates $\kappa$. (a) NRMSE as a function of $Delay$, with no feedback or 4 feedbacks, for $\kappa=10$. (b)(c) The actual (red) and target (blue) outputs for $Delay=200$ and $Delay=2$ with $\kappa=10$, corresponding to the points marked by letters “b” and “c” in panel (a), respectively. (d) NRMSE as a function of $\kappa$, with no feedback or 4 feedbacks, for $Delay=20$. (e)(f) The actual (red) and target (blue) outputs for $\kappa=10^{5}$ and $\kappa=10$ with $Delay=20$, corresponding to the points marked by letters “e” and “f” in panel (d), respectively. Parameters: $\omega_{1}=20$, $g_{1}=30$, $\omega_{c}=40$, $\epsilon=20$.
The impacts of $Delay$ and the decay rate $\kappa$ are shown in Fig. 4. In the Mackey-Glass task, a longer $Delay$ necessitates a stronger memory for the reservoir to “remember” more past information from the input in order to forecast the future output, making the task increasingly challenging. The decay rate $\kappa$ characterizes the loss of quantum memory caused by quantum dissipation, as define in the stochastic master equation introduced in the “Methods” section. A larger $\kappa$ tends to cause the reservoir to “forget” input information more quickly due to photon leakage and atomic spontaneous emission. However, a larger $\kappa$ is beneficial for stronger measurements with reduced uncertainties in readouts. Therefore, selecting a balanced $\kappa$ could be crucial in practice.
@@
Figure 6: Performance testing for the sine-square waveform classification task with various reservoir scales. (a)(b) Tested NRMSE plotted against the number of atoms (or neurons), using all available readout channels, 0, 22, or 44 feedback channels, and regular linear regression (solid lines) or polynomial regression (dashed lines), with a comparison to CRC employing classical ESN (red solid line), where the parameters $\omega_{i}$ and $g_{i}$ align with those in Fig. 3(a)(b). (c)(d) Actual (red) and target (blue) outputs from QRC utilizing one atom with no feedback with linear regression, and three atoms with 44 feedbacks with polynomial regression, corresponding to the points marked by “c” and “d” in panel (b), respectively. Parameters: $\omega_{c}=40$, $\kappa=10$, $\epsilon=20$, $\omega_{ss}=10$, and $N_{ss}=16$.
@@ II.3 Sine-square Waveform Classification Task
The objective of the classification task is to determine whether each input data point belongs to a sine or a square waveform. The time-dependent input, $f_{k}$, comprises $110$ randomly generated sine and square waveforms. Among these, $10$ waveforms are allocated for memory fading, $50$ for training, and $50$ for testing. Each waveform is discretized into $N_{ss}$ points, leading to a time step of $\Delta t=2\pi/\left(N_{ss}\omega_{ss}\right)$, where $\omega_{ss}$ denotes the oscillation frequency of the input. Figure 5(a) depicts these discretized time points (red dots) across the first $10$ waveforms during the testing phase. The corresponding input modified by feedback is plotted in Fig. 5(b). The target output, $\bar{y}_{k}$, aimed at classifying the input signal, is set to 0 if the input point belongs to a square waveform, and $1$ if it belongs to a sine waveform. Compared to the Mackey-Glass task, the sine-square task requires shorter-term memory, since the extremal points require memory to be distinguished–input values of 0 and $2$ can correspond to either the sine or the square wave.
@@
The sine-square waveform classification is a nonlinearity task that requires the reservoir to process a linearly inseparable input dataset containing abrupt changes. In a linear, closed quantum system, the smoothly evolving expectation values in readouts, $\left\langle c\right\rangle\propto\exp\left(-i\omega_{c}t\right)$ and $\left\langle\sigma_{i}\right\rangle\propto\exp\left(-i\omega_{i}t\right)$, are unable to capture these high-frequency, abrupt shifts. The introduction of nonlinearity, originating from the coherent laser driving and decay in the open quantum system, enhances the reservoir’s capability to promptly respond to the abrupt changes in the input. This is evidenced by the distinct readout measurements in Fig. 5(c)-(f), which correspond directly to the sine and square input waveforms in Fig. 5(a). The various feature shapes observed in the readout measurements highlight the reservoir’s capability to rapidly respond to the sudden shifts in input.
@@
The performance enhancements resulted from the increased number of atoms, feedback mechanism, and polynomial regression, along with the comparison between QRC and CRC, are shown in Figs. 6(a)(b). It is observed that the performance associated with polynomial regression, represented by the dashed lines, tends to saturate at $4$ atoms. This performance saturation suggests that the maximum performance achievable with the current sample size, associated with $N_{ss}=16$, has been attained. The enhanced performance is illustrated in Fig. 6(d), where discrepancies between the actual and target outputs primarily occur at the first point in each waveform. These discrepancies characterize the short-time responses of the quantum reservoir to the abrupt shifts in the input waveform. The dependence of QRC performance on various samples sizes $N_{ss}$ and the influence of abrupt shifts in the input are detailed in the “Supplementary Information”.
@@ III Discussion
We have introduced a paradigm for minimalistic quantum reservoirs that enhances quantum memory and nonlinear data processing capabilities. This approach incorporates a feedback mechanism that modifies the inputs based on observable readouts. Specifically, the inputs are encoded into the intensity of the cavity-driving laser, enabling a practical implementation of feedback by adjusting the laser’s intensity. In a previous work, a Hamiltonian configuration strategy was proposed to enhance QRC performance by using feedback from the outputs to modify the reservoir dynamics [Xia2023]. This approach is based on controlling internal parameters of the quantum reservoir, analogous to adjusting the atom-photon coupling strength $g_{i}$ and the detuning $\omega_{i}$ in our system. In contrast, our approach does not involve adjusting these internal parameters. Instead, the parameter $V_{n_{f}}$ is treated as an external parameter that lies outside the quantum reservoir hardware. The primary purpose of our external feedback is to provide an enhanced computational complexity by easily adjusting the weights $V_{n_{f}}$ through mostly software methods. This purpose is consistent with the original QRC paradigm, which does not require optimization of internal parameters.
To further augment the system, we employ polynomial regression to append quadratic combinations of observable expectations to the readouts, introducing an additional layer of complexity and capability. Similar nonlinear combinations of measured expectations have been leveraged in previous proposals based on qubit networks [CNY20, chen2019learning]. However, in these qubit-based systems, incorporating second- and third-order polynomial terms into the readouts has not led to significant improvements in NRMSE [chen2019learning]. In contrast, our photon-qubit coupled system benefits from the high-dimensional photon subspace, which enhances the reservoir’s expressive capacity. This potential is effectively harnessed through polynomial regression.
@@
These novel components significantly enhance the performance of QRC for small reservoirs containing just one to five atoms. To obtain the readouts in practice, we propose the use of continuous quantum measurements via homodyne detection of cavity quadratures and atomic spins. This approach allows for the simultaneous measurement of non-commuting observables [Wei2008, Ruskov2010, Hacohen2016, Ochoa2018], avoiding the need for tomography and making it more practical than the measurement of probability distributions in quantum basis states suggested in prior works [Fujii2017, Ghosh2019, Dudas2023, Angelatos2021].
@@
Our proposed quantum optical reservoir offers convenient scalability compared to reservoirs built upon quantum networks like the Ising [Fujii2017] and Fermi-Hubbard [Ghosh2019] models. This advantage arises from the practical coupling between atoms and a single-mode cavity field in our system. Specifically, to scale up the reservoir, we simply need to add more atoms into the cavity. These atoms naturally couple with the cavity field, which in turn automatically couples them with the rest of the atoms in the reservoir.
@@
Moreover, the number of quantum basis states in our reservoir scales exponentially with the number of atoms, following $2^{N_{\text{atom}}}$. This leads to faster growth compared to reservoirs that are not based on quantum networks. In systems such as a single Kerr nonlinear oscillator [Govia2021] or two coupled linear oscillators [Dudas2023], the number of basis states increases linearly with the energy levels involved in the dynamics. The exponential scaling of our reservoir has been crucial in demonstrating the advantages of QRC over CRC, as shown in Figs 3(a)(b) and 6(a)(b). To compare with CRC using the same number of neurons, Fig. 3(a) shows that QRC with 3 atoms significantly improves performance by reducing the NRMSE by $51\%$. For a 3-atom QRC, the introduction of a feedback mechanism with 4 feedback channels reduces the NRMSE by $42\%$, while polynomial regression alone contributes a $32\%$ reduction. When both the feedback mechanism and polynomial regression are applied together, the NRMSE reduction increases to $65\%$, as illustrated in Fig. 3(a).
The cooperativity serves as a key metric for assessing the feasibility of the proposed QRC scheme on realistic experimental platforms. It is a dimensionless parameter defined as $g_{i}^{2}/\kappa_{c}\kappa_{i}$, where $g_{i}$ denotes the atom-photon coupling strength, $\kappa_{c}$ is the cavity photon loss rate, and $\kappa_{i}$ is the atomic spontaneous emission rate. In our simulations, we use typical values of $g_{i}=10$ and $\kappa_{c}=\kappa_{i}=5$, resulting in a single-atom cooperativity of approximately $4$. Given that many experimental platforms operate with cooperativity values ranging from $1$ to $10$ [Reitz2022], our simulated system is within realistic experimental conditions.
@@ IV.1 Continuous quantum measurement for observable readouts
Continuous quantum measurements are simulated for both the cavity field and individual atoms. For the cavity field, homodyne detection of two orthogonal quadratures involves splitting the system’s output beam into two using a beam-splitter, followed by homodyning each beam with the same local oscillator, differing by a phase shift of $\pi/2$ [Wiseman2010]. Similarly, for atoms, homodyne detection is performed for spontaneous emissions [Wiseman2001]. These measurements can be conducted concurrently [Wei2008, Ruskov2010, Hacohen2016, Ochoa2018]. Consequently, the continuous measurement process is described by the stochastic master equation [VPB79, VPB88, VPB91a, HC93, Wiseman2010, BvHJ07, Nurdin14]
@@
$+2\mathcal{D}\left[\sqrt{\kappa_{c}}c\right]\rho_{J}dt+2\underset{i}{\sum}\mathcal{D}\left[\sqrt{\kappa_{i}}\sigma_{i}\right]\rho_{J}dt$
@@
$+\underset{i}{\sum}\left(dW_{x,i}\mathcal{H}\left[\sqrt{\kappa_{i}}\sigma_{i}\right]+dW_{y,i}\mathcal{H}\left[i\sqrt{\kappa_{i}}\sigma_{i}\right]\right)\rho_{J},$
@@
for any stochastic collapse operator $a$. The continuous quantum measurements of the observables $Q$, $P$, $\sigma_{x,i}$, and $\sigma_{y,i}$ are associated with the stochastic collapse operators $\sqrt{\kappa_{c}}c$, $i\sqrt{\kappa_{c}}c$, $\sqrt{\kappa_{i}}\sigma_{i}$, and $i\sqrt{\kappa_{i}}\sigma_{i}$, respectively [Wiseman2010, BvHJ07, Nurdin14]. The randomness of the measurement records is taken into account by the Wiener increments, $dW_{Q\left(P\right)}$ and $dW_{x\left(y\right),i}$, each of which selects a random number from a Gaussian probability distribution with a width of $dt$. The efficiencies of various detection channels are incorporated into the Wiener increments. Each measurement detects the continuous currents in cavity and atom channels with noises, with the measurement records given by $\left\langle Q\left(P\right)\right\rangle_{J}+dW_{Q\left(P\right)}/dt$ and $\left\langle\sigma_{x\left(y\right),i}\right\rangle_{J}+dW_{x\left(y\right),i}/dt$, where the expectation values are computed using $\rho_{J}$ from Eq. (11). The “smesolve” method in the Python QuTiP library is employed to solve the stochastic master equation (11).
@@
$\frac{d\rho}{dt}=-i\left[H_{0}+H_{1}\left(t\right),\rho\right]+2\mathcal{D}\left[\sqrt{\kappa_{c}}c\right]\rho+2\underset{i}{\sum}\mathcal{D}\left[\sqrt{\kappa_{i}}\sigma_{i}\right]\rho,$ (14)
@@
accompanied by the averaged measurement records $\left\langle Q\left(P\right)\right\rangle$ and $\left\langle\sigma_{x\left(y\right),i}\right\rangle$, where the expectation values are computed using $\rho$ from Eq. (14). To maintain a constant total decay rate $\kappa$ as $N_{atom}$ increases, the decay rate of the cavity or each atom is assumed to be $\kappa_{c\left(i\right)}=\kappa/\left(2N_{atom}+2\right)$. A higher $\kappa$ reduces the uncertainty in the readout measurement, commonly referred to a strong measurement, while a lower $\kappa$ corresponds to a weak measurement [Fuchs2001]. The influence of the magnitude of $\kappa$ on QRC performance is demonstrated in Fig. 4. The “mesolve” method in the Python QuTiP library is employed to solve the deterministic master equation (14).
Figure 7: Testing results for the Mackey-Glass task within the framework of continuous quantum measurement, where each readout is obtained by averaging measurement records over multiple measurement trajectories, with each measurement trajectory running through the whole time domain. (a) NRMSE as a function of the number of measurement trajectories for a single-atom QRC with $g_{1}=30$ and $\omega_{1}=20$. Solid lines: averaged results from a finite number of trajectories simulated with stochastic master equation. Dashed lines: asymptotic results simulated with deterministic master equation. Green lines: 44 feedbacks. Black lines: No feedback. (b)(c) The actual (red) and target (blue) outputs for $10$ and $10000$ trajectories with 44 feedbacks, corresponding to the points labeled “b” and “c” in panel (a), respectively. Parameters: $Delay=20$, $\kappa=10$, $\omega_{c}=40$, $\epsilon=20$.
To eliminate the need for many copies of quantum systems and many repetitions of measurement to estimate observable expectations at each time step, we propose an online protocol based on continuous quantum measurement. This method requires only a single quantum system and generates multiple quantum trajectories, each evolving over the entire physical time domain, thereby avoiding destructive measurements. These trajectories are implemented sequentially on the same quantum system, with a memory-fading time interval introduced between each pair of successive trajectories. The fading memory feature of our QRC scheme eliminates the need to reset the initial state before the start of each trajectory.
For the QRC scheme without feedback, a continuous measurement trajectory is simulated by evolving the stochastic master equation in Eq. (11) over the entire time domain from $t=0$ to $t=L\Delta t$, where $L$ is the maximum value of $k$ in the discretized time $t=k\Delta t$. The test phase of QRC is finalized at $t=L\Delta t$. The required number of experimental repetitions is equal to the number of measurement trajectories, denoted as $N_{Trajectories}$. To simulate the stochastic nature of measurement noise, the stochastic master equation in Eq. (11) is numerically evolved multiple times, each with independent realizations of the random Wiener increments $dW_{Q\left(P\right)}$ and $dW_{x\left(y\right),i}$. Each full evolution corresponds to a single measurement trajectory. The observable readouts, $x_{kn}$, are then obtained by averaging the measurement records across all such trajectories. The effect of using a finite number of measurement trajectories is illustrated by the black solid line in Fig. 7(a) for a single-atom QRC without feedback. The idealized case, based on observables averaged over an infinite number of trajectories, is simulated using the deterministic master equation (14), with the result plotted by the flat black dashed line in Fig. 7(a). As shown in Fig. 7(a), increasing the number of trajectories leads to improved QRC performance, with the result for $N_{Trajectories}=10000$ approaching the limit obtained from an infinite number of trajectories. In the “Results” section, we use the deterministic master equation method to characterize QRC performance.
For the QRC scheme enhanced with feedback, our continuous measurement approach also allows each measurement trajectory to span the entire time domain from $t=0$ to $t=L\Delta t$ without destroying the quantum reservoir. This eliminates the need for a large number of measurement repetitions to estimate observable expectations at each time step. In the feedback scheme, the required number of experimental repetitions remains equal to the number of measurement trajectories. For each trajectory, the system evolves from $t_{k}$ to $t_{k+1}$ according to the stochastic master equation (11), during which the modified input and density operator are updated from $\widetilde{f}_{k}$ and $\rho_{J}\left(t_{k}\right)$ to $\widetilde{f}_{k+1}$ and $\rho_{J}\left(t_{k+1}\right)$, respectively. At the time $t_{k+1}$, the evolution is uninterrupted as we only take one non-destructive measurement shot from the current trajectory. To ensure the stability of the feedback mechanism, the modified input $\widetilde{f}_{k}$ for the $M$-th trajectory is determined by
$\widetilde{f}_{k}=f_{k}+\underset{n_{f}}{\sum}\overline{x}_{kn_{f}}^{M}V_{n_{f}},$ (15)
where $\overline{x}_{kn_{f}}^{M}$ is the averaged readouts from all previous measurement trajectories indexed from $1$ to (M1)(M-1), which is given by
$\overline{x}_{kn_{f}}^{M}=\frac{1}{M-1}\mathop{\sum_{m=1}^{M-1}}x_{kn_{f}}^{m},$ (16)
with $m$ denoting the index of a previous trajectory and $x_{kn_{f}}^{m}$ the readout from the $m$-th trajectory. Note that the readouts from the current ($M$-th) trajectory, denoted as $x_{kn_{f}}^{M}$, are determined by the solution $\rho_{J}\left(t_{k}\right)$ from the stochastic master equation of the current trajectory. The idealized case for feedback QRC is simulated using the deterministic master equation (14), using the feedback modified input governed by Eq. (1). In the “Results” section, we use the deterministic master equation to evaluate the performance of QRC enhanced with feedback.
The performance of feedback QRC with a finite number of measurement trajectories is illustrated by the green solid line in Fig. 7(a). The averaged measurement result asymptotically approaches the NRMSE obtained using the deterministic master equation. Notably, measurement noise has a more pronounced effect on the feedback scheme compared to the scheme without feedback, as evidenced by the larger gap between the green solid and dashed lines than that between the black solid and dashed lines in Fig. 7(a). This observation is consistent with the fact that measurement noise affects not only the readouts from the quantum reservoir but also the inputs modified via feedback based on those readouts. Strikingly, as shown in Fig. 7(a), even in the presence of measurement noise, the feedback-enhanced QRC outperforms the idealized QRC without feedback when the number of trajectories exceeds $1000$.